A lorry and a car travel uniformly along expressway from city A to city B . Thw lorry leaves at 915 amd arrives at 1315. The car leaves at 1000 and arrive at 1245. At what time did the car overtake the lorry.?
Answer by fcabanski(1391)

d = r * t (Distance is rate times time).


Call the lorry's speed L and the car's speed C.


The distance they travel is the same, so 4L = 2.75C


The distance the car must make up in some time (t) is the time of head start the lorry had, times the lorry's speed, which is: .75L and that equals the car's speed times the time (t) so .75L = Ct


Use the first equation to find L in terms of C. 4L=2.75C so L = 2.75/4 * C = .6875C.


Substitute that into the second equation:
.75(.6875C) = Ct so .515625C = Ct and .515625 = t = approx 30.9 minutes.


Contact me for private tutoring: fcabanski@hotmail.com